OpenStudy (solomonzelman):

An inverse trig-function, inside a (reg) trig-function. (Examples - I will be adding examples from time to time)

OpenStudy (solomonzelman):

\({\bbox[5pt, #99ffcc ,border:2px solid black ]{ \bf Example~~\text{#}1 }}\) \(\color{#000000 }{ \displaystyle \sin\left(\arccos \sqrt{1/4}\right) }\) We know that: \(\color{#000000 }{ \displaystyle \sin^2\theta+\cos^2\theta=1 }\) \(\color{#000000 }{ \displaystyle \sin^2\theta=1-\cos^2\theta }\) \(\color{#000000 }{ \displaystyle \sqrt{\sin^2\theta}=\sqrt{1-\cos^2\theta} }\) \(\color{#0000ff }{ \displaystyle \sin\theta=\sqrt{1-\cos^2\theta} }\) In our case, \(\theta\) is \(\color{#000000 }{ \displaystyle\left(\arccos \sqrt{1/4}\right) }\). So will apply this, \(\color{#000000 }{ \displaystyle \sqrt{1-\cos^2\left(\arccos \sqrt{1/4}\right)} }\) Knowing that \(\color{#000000 }{ \displaystyle \cos\left(\arccos x\right)=x }\) thus, \(\color{#000000 }{ \displaystyle \cos^2\left(\arccos x\right)=x^2 }\) and so is any, \(\color{#000000 }{ \displaystyle \cos^n\left(\arccos x\right)=x^n }\) So, we have, \(\color{#000000 }{ \displaystyle \sqrt{1-\left[ \sqrt{1/4}\right]^2} }\) \(\color{#000000 }{ \displaystyle \sqrt{1-(1/4)} }\) \(\color{#000000 }{ \displaystyle \sqrt{3/4} =\sqrt{3}/2}\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sec\left(\arccos \frac{1}{2}\right) }\) since \(\sec\theta = 1/\cos\theta\), \(\color{#000000 }{ \displaystyle \frac{1}{\cos\left(\arccos \frac{1}{2}\right) } }\) We apply the same rule, \(\cos(\arccos x)=x\), and we therefore get: \(\color{#000000 }{ \displaystyle \frac{1}{~\dfrac{1}{2}~} \Rightarrow 1\div \frac{1}{2}\Rightarrow1\times \frac{2}{1}\Rightarrow\color{blue}{2} }\) (I will start coloring the answer in blue for convinience)

OpenStudy (alexandervonhumboldt2):

\(\Huge\color{white}{yay}\)

OpenStudy (alexandervonhumboldt2):

*claps* *sending appalade* xD

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \tan(\arcsin 0.5 ) }\) since \(\tan \theta = \sin \theta / \cos\theta \) \(\color{#000000 }{ \displaystyle \frac{\sin(\arcsin 0.5 )}{\cos(\arcsin 0.5 )} }\) the top comes out right away... Another important rule (of same nature as this rule by the cosine function) \(\sin(\arcsin x )=x\) \(\color{#000000 }{ \displaystyle \frac{0.5}{\cos(\arcsin 0.5 )} }\) Knowing the Pythegorean identity, \(\color{#000000 }{ \displaystyle \sin^2\theta +\cos^2\theta =1 }\) we can obtain that \(\color{#000000 }{ \displaystyle \cos^2\theta =1 - \sin^2\theta }\) \(\color{#000000 }{ \displaystyle \cos \theta =\sqrt{1 - \sin^2\theta} }\) And we can apply the above to our denominator \(\color{#000000 }{ \displaystyle \frac{0.5}{\sqrt{1 - \sin^2(\arcsin 0.5 )}} }\) Same as by the cosine, \(\cos^n(\arccos x)=x^n\) you have the same by the sine function, \(\sin (\arcsin x)=x \) as well as \(\sin^n(\arcsin x)=x^n\) (same as by the cosine) ... And we get, \(\color{#000000 }{ \displaystyle \frac{0.5}{\sqrt{1 - \sin^2(\arcsin 0.5 )}} }\) \(\color{#000000 }{ \displaystyle \frac{0.5}{\sqrt{1 - (0.5)^2}} }\) \(\color{#000000 }{ \displaystyle \frac{0.5}{\sqrt{1 - 0.25}} }\) \(\color{#000000 }{ \displaystyle \frac{0.5}{\sqrt{ 0.75}} }\) and then the rest is a matter of algebra, \(\color{#000000 }{ \displaystyle \frac{0.5}{\sqrt{ 75/100}} = \frac{\dfrac{1}{2}}{\dfrac{\sqrt{3}}{2}}=1/\sqrt{3} =\color{blue}{\sqrt{3}/3} }\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sin(\arctan e) }\) We need to express sine in terms of tangent, \(\tan \theta=\sin\theta/\cos \theta \) \(\sin\theta=\tan \theta\cos \theta \) \(\sin\theta=\tan \theta\left(\dfrac{1}{\sec \theta }\right) \) \(\color{royalblue}{\sec^2\theta=\tan^2\theta+1}\) \(\color{royalblue}{\sec\theta=\sqrt{\tan^2\theta+1}}\) \(\sin\theta=\dfrac{\tan \theta}{\sqrt{1-\tan^2\theta } } \) therefore, \(\color{#000000 }{ \displaystyle \sin(\arctan e)=\dfrac{\tan (\arctan e)}{\sqrt{1-\tan^2(\arctan e)~ } {\huge\color{white}{|}}} }\) the rules of tangent are same by cosine and sine (and by every trig function) That is, that: \(\tan(\arctan x)=x \), \(\tan^2(\arctan x)=x^2 \) and any, \(\tan^n(\arctan x)=x^n \) Therefore, we get, \(\color{#000000 }{\sin(\arctan e)=\dfrac{e}{\sqrt{1-e^2 }} }\)